OpenStudy (samigupta8):

five persons a,b,c,d,e are pulling a cart of mass 100 kg on a smooth surface and cart is moving with acceleration 3m/s^2 in east direction.when person a stops pulling it moves with acceleration 1m/s^2 in west direction.when person b stops pulling it moves with acceleration 24m/s^2 in north direction.the magnitude of acceleration of the cart when only a and b pull the cart keeping their directions same as the old direction is

OpenStudy (anonymous):

here a+b+c+d+e= 300 i Newton b+c+d+e= -100 i N from here a= 400 i N now a+c+d+e = 2400 j N so b= (300 i-2400 j) N now if only and b act on the body the force on the body is a+b= (700 i - 2400 j) N here old direction is east then acceleration in east direction is=700/100= 7 m/s2 is it correct??

OpenStudy (samigupta8):

no it's 25 m/s^2

OpenStudy (samigupta8):

we cn take the magnitude of a=\[\sqrt{700^{2}+2400^{2}}\]

OpenStudy (samigupta8):

sorry it's magnitude of a+b

OpenStudy (samigupta8):

it will come out to be a+b=2500m/s^2

OpenStudy (samigupta8):

bt i m not getting one thing why are v dividing it by 100 to get 25

OpenStudy (anonymous):

a=f/m